### v1

# Dahyeon Kye, Changhyun Roh, Sukhun Ko, Chanho Eom, and Jihyong Oh https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation

Abstract
Video Frame Interpolation (VFI) is a fundamental Low-Level Vision (LLV) task that synthesizes intermediate frames between existing ones while maintaining spatial and temporal coherence. VFI techniques have evolved from classical motion compensation-based approach to deep learning-based approach, including kernel-, flow-, hybrid-, phase-, GAN-, Transformer-, Mamba-, and more recently diffusion model-based approach. We introduce AceVFI, the most comprehensive survey on VFI to date, covering over 250+ papers across these approaches. We systematically organize and describe VFI methodologies, detailing the core principles, design assumptions, and technical characteristics of each approach. We categorize the learning paradigm of VFI methods namely, Center-Time Frame Interpolation (CTFI) and Arbitrary-Time Frame Interpolation (ATFI). We analyze key challenges of VFI such as large motion, occlusion, lighting variation, and non-linear motion. In addition, we review standard datasets, loss functions, evaluation metrics. We examine applications of VFI including event-based, cartoon, medical image VFI and joint VFI with other LLV tasks. We conclude by outlining promising future research directions to support continued progress in the field. This survey aims to serve as a unified reference for both newcomers and experts seeking a deep understanding of modern VFI landscapes. We maintain an up-to-date project page: https://github.com/CMLab-Korea/Awesome-Video-Frame-Interpolation.
Index Terms:
Video Frame Interpolation, Generative Inbetweening, Video Generation, Low-Level Vision
publicationid: pubid:
1 Introduction
Video Frame Interpolation (VFI) aims to increase the temporal resolution (i.e., frame rate) of a video sequence by synthesizing one or more intermediate frames between given consecutive frames. This task serves a broad range of applications, including novel view synthesis [1, 2, 3, 4], slow-motion generation [5, 6, 7, 8, 9, 10], video compression [11, 12, 13, 14], video prediction [15, 16, 13, 17], and diverse generation tasks such as co-speech reenactment [18], human motion synthesis [19], and facial animation [20]. A key advantage of VFI lies in its ability to synthesize perceptually smooth and temporally coherent motion, aligning well with the temporal characteristics of the human visual system (HVS). High-frame-rate (HFR) content reduces artifacts such as motion blur and judder [21, 22], thereby enhancing the visual quality in high-resolution (HR) and immersive media. This makes VFI particularly valuable in latency-sensitive and fidelity-critical scenarios such as sports broadcasting, interactive gaming, and virtual reality. Finally, in streaming pipelines, VFI also enables bandwidth-efficient video transmission by reconstructing intermediate frames locally, reducing the need to transmit full frame sequences [21].
Figure 1: General process of VFI. Given two input frames $I_{0}$ and $I_{1}$, the VFI model $\mathcal{F}$ synthesizes one or more intermediate frames. ×$\times n$ interpolation denotes synthesizing $n{-}1$ intermediate frames to increase the frame rate by a factor of $n$.
Formally, given two frames $I_{0}$ and $I_{1}$, a VFI model $\mathcal{F}$ estimates the interpolated frame $\hat{I}_{t}$ at time $t\in(0,1)$:
$\hat{I}_{t}=\mathcal{F}(I_{0},I_{1},t).$ (1)
As shown in Fig. 1, interpolating $n{-}1$ frames between each input pair increases the frame rate by a factor of $n$. For example, generating seven intermediate frames per interval transforms a 30fps video into 240fps.
1.1 General Pipeline of VFI
[…]
Figure 2: General pipeline of VFI. Given two input frames $I_{0}$ and $I_{1}$, the feature representations $F_{0}$ and $F_{1}$ are extracted, which are then aligned to the target time $t$ using the estimated motion. The temporal alignment can be performed on either input RGB pixels or features, resulting in $\hat{I}_{0\rightarrow t},\hat{I}_{1\rightarrow t}$ or $\hat{F}_{0\rightarrow t},\hat{F}_{1\rightarrow t}$. Finally, a Frame Synthesis module blends the aligned inputs to produce the interpolated frame $\hat{I}_{t}$. This pipeline highlights the four core stages of VFI: Feature Extraction, Motion Estimation, Temporal Alignment, and Frame Synthesis.
As shown in Fig. 2, the general VFI pipeline consists of four stages. (i) Feature Extraction: Input frames $I_{0}$ and $I_{1}$ are first passed through a feature extraction network [23, 24, 25] to obtain deep features $F_{0}$ and $F_{1}$. These features encode spatial and semantic information suitable for subsequent motion reasoning [26]. (ii) Motion Estimation: The temporal correspondence, commonly referred to as motion, is estimated. Motion estimation is performed either explicitly (e.g., optical flow [27]) or implicitly via kernels [28, 29, 16, 30, 31, 32, 33, 34, 35, 36, 37, 8, 38, 39, 40, 41, 42, 43, 44, 45, 46, 47], phase [48, 49], attention maps [50, 51, 52, 53, 54, 55, 56, 57], or cost volumes [35, 58, 59, 53]. (iii) Temporal Alignment: The estimated motion is used to temporally align the input pixels or features to the target time $t$, resulting in $\hat{I}_{0\rightarrow t}$, $\hat{I}_{1\rightarrow t}$ or $\hat{F}_{0\rightarrow t}$, $\hat{F}_{1\rightarrow t}$. There are four types of alignment strategies. Kernel-based alignment (Fig. 3 (a)) aggregates local or non-local information from the inputs using learned, spatially-adaptive kernels. These kernels implicitly encode motion by adapting their spatial weights based on local context, allowing motion-aware alignment without explicit flow estimation. Flow-based alignment (Fig. 3 (b)) warps inputs guided by the estimated flow. Forward warping [60] maps source pixels (i.e., input pixels) to their estimated locations in the target frame. Backward warping [61] samples from the source based on coordinates in the target frame, effectively pulling information from the source toward the desired time. Attention-based alignment (Fig. 3 (c)) replaces explicit geometric warping with attention-weighted aggregation [51, 53]. By computing soft correspondences between elements across input frames, this approach can adaptively focus on semantically relevant regions and align contents even across large spatial-temporal gaps. Cost volume-based alignment (Fig. 3 (d)) constructs dense similarity volumes between feature maps, enabling precise correspondence modeling across space and time. (iv) Frame Synthesis: Finally, the aligned inputs are blended to synthesize the final interpolated frame using either simple averaging, weighted blending or synthesis networks [62].
Figure 3: Temporal alignment process. The input frames $(I_{0},I_{1})$ or their features $(F_{0},F_{1})$ are temporally aligned toward a target time $t$. (a) kernel-based alignment using spatially-adaptive filters, (b) flow-based alignment guided by optical flow, (c) attention-based alignment using attention-weighted correspondences, and (d) cost volume-based alignment through pixel- or feature-level similarity estimation.
{forest}
[…]
Combined (Sec. 2.2.3), color=colorSec2!95, fill=colorSec2!25, text=black] [Phase-based (Sec. 2.2.4), color=colorSec2, fill=colorSec2!25, text=black] [GAN-based (Sec. 2.2.5), color=colorSec2, fill=colorSec2!25, text=black] [Transformer-based (Sec. 2.2.6), color=colorSec2, fill=colorSec2!25, text=black] [Mamba-based (Sec. 2.2.7), color=colorSec2, fill=colorSec2!25, text=black] ] [Diffusion-based (Sec. 2.3), color=colorSec2, fill=colorSec2!50, text=black] ] [III. Learning
Paradigm (Sec. 3), color=colorSec3, fill=colorSec3, text=black [Center-Time Frame Interpolation
(CTFI) (Sec. 3.1), color=colorSec3, fill=colorSec3!50, text=black] [Arbitrary-Time Frame Interpolation
(ATFI) (Sec. 3.2), color=colorSec3, fill=colorSec3!50, text=black] [Training Strategy (Sec. 3.3), color=colorSec3!95, fill=colorSec3!30, text=black [CTFI Training Strategy
(CTFI-TS) (Sec. 3.3.1), color=colorSec3, fill=colorSec3!25, text=black] [ATFI Training Strategy
(ATFI-TS) (Sec. 3.3.2), color=colorSec3, fill=colorSec3!25, text=black] ] [Loss Functions (Sec. 3.4), color=colorSec3, fill=colorSec3!50, text=black [Reconstruction Loss (Sec. 3.4.1), color=colorSec3, fill=colorSec3!25, text=black] [Perceptual Loss (Sec. 3.4.2), color=colorSec3, fill=colorSec3!25, text=black] [Adversarial Loss (Sec. 3.4.3), color=colorSec3, fill=colorSec3!25, text=black] [Flow Loss (Sec. 3.4.4), color=colorSec3, fill=colorSec3!25, text=black] ] ] [IV. VFI Challenges (Sec. 4), color=colorSec4, fill=colorSec4, text=black [Large Motion (Sec. 4.1), color=colorSec4, fill=colorSec4!50, text=black] [Occlusion (Sec. 4.2), color=colorSec4, fill=colorSec4!50, text=black] [Lighting Variation (Sec. 4.3), color=colorSec4, fill=colorSec4!50, text=black] [Non-linear Motion (Sec. 4.4), color=colorSec4, fill=colorSec4!50, text=black] ] [V. Datasets and
Evaluation Metrics (Sec. 5), color=colorSec5, fill=colorSec5, text=black [Datasets (Sec. 5.1), color=colorSec5, fill=colorSec5!50, text=black [Triplet Datasets (Sec. 5.1.1), color=colorSec5, fill=colorSec5!25, text=black] [Multi-frame Datasets (Sec. 5.1.2), color=colorSec5, fill=colorSec5!25, text=black] ] [Data Augmentation (Sec. 5.2), color=colorSec5, fill=colorSec5!50, text=black] [Evaluation Metrics (Sec. 5.3), color=colorSec5, fill=colorSec5!50, text=black [Image-level Metrics (Sec. 5.3.1), color=colorSec5, fill=colorSec5!25, text=black] [Perceptual Metrics (Sec. 5.3.2), color=colorSec5, fill=colorSec5!25, text=black] [Video-level Metrics (Sec. 5.3.3), color=colorSec5, fill=colorSec5!25, text=black] ] ] [VI. Applications (Sec. 6), color=colorSec6, fill=colorSec6, text=black [Event-based VFI (Sec. 6.1), color=colorSec6, fill=colorSec6!50, text=black] [Cartoon VFI (Sec. 6.2), color=colorSec6, fill=colorSec6!50, text=black] [Medical Image VFI (Sec. 6.3), color=colorSec6, fill=colorSec6!50, text=black] [Joint Task VFI (Sec. 6.4), color=colorSec6, fill=colorSec6!50, text=black] ] [VII. Future Research
Directions (Sec. 7), color=colorSec7, fill=colorSec7, text=black [Video Streaming Service (Sec. 7.1), color=colorSec7, fill=colorSec7!50, text=black] [All-in-One
LLV Video Restoration (Sec. 7.2), color=colorSec7, fill=colorSec7!50, text=black] [3D and 4D
Scene Understanding (Sec. 7.3), color=colorSec7, fill=colorSec7!50, text=black] ] ]
Figure 4: Overview of the survey structure. The figure summarizes the hierarchical organization of the survey, including methodological categories (Sec. 2), learning paradigms (Sec. 3), key challenges (Sec. 4), datasets and evaluation metrics (Sec. 5), applications (Sec. 6), and future research directions (Sec. 7).
1.2 Methodology Overview
VFI methodologies can be broadly classified into three major categories: motion compensation-based [63, 64, 65, 66, 67, 68, 69, 70, 71, 72], deep learning-based [61, 16, 73, 7, 74, 75, 76, 62, 60, 77, 78, 79, 80, 81, 58, 59, 82, 83, 84, 55, 85, 54, 53, 86, 87, 88, 89, 90, 91, 92, 28, 29, 32, 31, 50, 34, 39, 40, 41, 42, 43, 44, 45, 46, 30, 33, 35, 36, 37, 38, 93, 94, 95, 96, 74, 97, 48, 49, 51, 52, 56, 57, 98, 99, 100], and diffusion models (DMs)-based approach [101, 96, 102, 103, 104, 105, 106, 107, 108, 109, 110, 111, 112, 113, 114, 115, 116, 117, 118, 119].
The motion compensation-based approach dominated the pre-deep-learning era, offering a straightforward two-stage strategy: estimating motion explicitly and warping frames accordingly. While effective under simple motion, its reliance on hand-crafted rules and block-based assumptions limits its ability to handle occlusions and complex, non-rigid dynamics.
With the advent of convolutional neural networks (CNNs) [120], the field shifted toward deep learning-based approach. This approach replaces heuristic pipelines with end-to-end frameworks that learn motion patterns and appearance features directly from data. As a result, they significantly improve robustness under diverse and challenging conditions. Further methodological details are discussed in Sec. 2.2.
More recently, DMs have been introduced as a generative perspective to VFI, framing it as a conditional denoising process rather than a deterministic prediction task. This has expanded the scope of VFI into the broader paradigm of Generative Inbetweening [112, 111], enabling uncertainty-aware interpolation and semantically diverse frame synthesis. This shift not only enhances robustness in ambiguous motion scenarios but also opens the door to multi-modal guidance (e.g., text, depth, or motion priors), redefining the role of VFI in creative and interactive video generation.
Overview. Fig. 4 shows the overall structure of this paper. Sec. 2 analyzes methodological taxonomies of VFI. Sec. 3 introduces and compares the two principal learning paradigms of VFI, and further examines their corresponding training strategies and loss functions. Sec. 4 discusses major challenges in VFI, along with how recent methods address them. Sec. 5 reviews common datasets and evaluation metrics. Sec. 6 explores applications of VFI across diverse domains. Finally, Sec. 7 presents future research directions of VFI.
2 Methodology
2.1 Motion Compensation-based
Before the advent of deep-learning, VFI was primarily tackled using Motion-Compensated Frame Interpolation (MCFI) [64, 67, 68, 70] or Frame Rate Up-Conversion (FRUC) [63, 65, 66, 69, 71, 72], which dominated the field from the late 1990s through the early 2000s. These approaches explicitly estimate motion, typically via block matching or global parametric models, and synthesize the intermediate frame by warping the input frames according to the estimated motion fields. A typical MCFI pipeline involves two steps: (i) block-based motion estimation and (ii) pixel-level warping for frame synthesis. In block-based estimation, the frame is partitioned into fixed-size rectangular blocks, assuming uniform motion within each region. While this formulation offers computational efficiency, it fails to capture non-rigid or object-specific motion, often resulting in artifacts such as holes (due to occlusions) and overlaps (due to many-to-one mappings). Rooted in classical video coding frameworks [121], MCFI methods emphasize speed and simplcity, but inherently lack the capacity to handle fine-grained, non-linear motion. To address these issues, several extensions are proposed, including multi-stage estimation [70], adaptive motion models [66], and occlusion-aware warping [72]. Intermediate frame synthesis is typically achieved via block-wise projection or forward guided by the estimated motion.
Despite their limited robustness under complex dynamics, MCFI and FRUC methods [63, 64, 65, 66, 67, 68, 69, 70, 71, 72] lay the conceptual foundation for modern VFI. This core principle, which involves explicit motion estimation followed by motion-guided warping, remains central to many modern learning-based models and is now enhanced with deep feature representations and end-to-end training. Importantly, classical motion-compensated strategies offer valuable insights into the inductive biases that shape modern VFI architectures. Concepts such as motion locality, piecewise rigidity, and spatial warping, which originated from block-based estimation, are implicitly retained in modern mechanisms including deformable convolutions [122, 32] and local attention [123]. Furthermore, the challenges encountered in this approach, such as occlusion handling and motion discontinuity, have directly influenced the design of occlusion-aware blending modules and bidirectional flow formulations in recent models. In this light, traditional motion models serve as both a historical foundation and a conceptual framework for the progressive development of VFI architectures.
2.2 Deep Learning-based
2.2.1 Kernel-based
[…]
Figure 5: Comparison of different convolution types. (a) Standard convolution samples at a fixed grid location $(x+k,y+l)$. (b) Deformable convolution introduces learnable offsets $(\alpha_{k,l},\beta_{k,l})$, enabling adaptive sampling at $(x+k+\alpha_{k,l},\ y+l+\beta_{k,l})$. (c) Dynamic convolution further generalizes this by predicting the kernel weights $W_{i}^{k,l}(x,y)$ dynamically for each output position, allowing for spatially-variant filtering.
Kernel-based VFI methods [28, 29, 16, 30, 31, 32, 33, 34, 35, 36, 37, 8, 38, 39, 40, 41, 42, 43, 44, 45, 46, 56, 47] synthesize intermediate frames by predicting spatially-adaptive convolutional kernels, which are subsequently applied to local patches extracted from the input frames. Motion information is implicitly encoded in the kernel weights, thereby enabling motion-aware pixel aggregation without explicit motion estimation. As shown in Fig. 5 (a), a standard kernel-based interpolation can be mathematically formulated as:
$\hat{I}(x,y)=\sum_{i=0}^{N-1}\sum_{k=0}^{R-1}\sum_{l=0}^{R-1}W_{k,l}I_{i}(x+k,y+l),$ (2)
where $N$ denotes the number of input frames, $R$ denotes the kernel size, and $W_{k,l}$ represents the learned kernel weight at offset $(k,l)$. This approach adopts a simple single-stage formulation that combines motion estimation and frame synthesis into a one-step process. AdaConv [28] utilizes a U-Net-like architecture [24] to predict spatially-varying 2D kernels for each output pixel. This enables local, pixel-wise motion-aware aggregation that can implicitly handle both alignment and occlusion [46]. SepConv [29] further reduces the computational overhead by decomposing the 2D kernel into separable 1D kernels:
$W=W_{v}*W_{h},$ (3)
where $W_{v}\in\mathbb{R}^{R\times 1}$ and $W_{h}\in\mathbb{R}^{1\times R}$ are vertical and horizontal 1D kernels respectively. The * denotes the outer product between the two 1D kernels, resulting in a full 2D kernel $W\in\mathbb{R}^{R\times R}$. This decomposition reduces the number of learnable parameters from $R^{2}$ to $2R$, while maintaining a comparable receptive field. Despite their conceptual simplicity, these methods are inherently limited in handling large displacements due to their fixed receptive fields [74]. Such constraints stem from the content-agnostic nature of CNNs, which uniformly apply learned filters across spatial locations [51]. While this weight-sharing inductive bias proves effective in recognition tasks, it becomes suboptimal in VFI, where fine-grained motion modeling is essential. To overcome this problem, deformable kernel-based methods [32, 34, 37, 8, 40, 42, 39, 44, 51, 56] introduce learnable offsets [122] as shown Fig. 5 (b), which allow sampling outside the regular convolution grid:
$\hat{I}_{(}x,y)=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{R-1}\sum_{l=0}^{R-1}W_{k,l}$
$\cdot I_{i}(x+k+\alpha_{k,l},\ y+l+\beta_{k,l}),$ (4)
where $(\alpha_{k,l},\beta_{k,l})$ are the learnable offsets. AdaCoF [36] estimates both kernel weights and sampling offsets for each output pixel, though it employs a fixed offset pattern, limiting its expressiveness under complex motion. To enhance spatial adaptivity further as shown in Fig. 5 (c), dynamic kernel-based methods [58, 35, 41, 44, 124] make location-dependent kernel weights:
$\hat{I}(x,y)=$ $\sum_{i=0}^{N-1}\sum_{k=0}^{R-1}\sum_{l=0}^{R-1}W_{k,l}(x,y)$
$\cdot I_{i}(x+k+\alpha_{k,l},\ y+l+\beta_{k,l}),$ (5)
where $W_{k,l}(x,y)$ denotes a dynamically predicted kernel at location $(x,y)$. Methods such as CDFI [41] and MSEConv [44] jointly learn spatially-varying weights and offsets, offering enhanced flexibility and improved interpolation accuracy.
Overall, kernel-based methods forego explicit motion supervision, instead leveraging learned spatial priors for synthesizing intermediate frames. As kernel prediction and convolution-based synthesis are tightly coupled, motion estimation and frame synthesis are implicitly fused into a single-stage. This implicit formulation offers robustness in noisy or uncertain motion settings and eliminates the dependency on accurate optical flow. However, these models often lack temporal generalizability, as the learned kernels are typically conditioned on a fixed interpolation time (e.g., $t=0.5$). Consequently, most kernel-based methods are constrained to CTFI (Sec. 3.1) and fail to generalize to ATFI (Sec. 3.2), limiting their applicability in real-world scenarios requiring temporal flexibility.
2.2.2 Flow-based
Flow-based methods [16, 73, 74, 75, 76, 62, 60, 77, 78, 79, 39, 80, 81, 58, 59, 82, 83, 84, 125, 55, 85, 54, 53, 87, 88, 89, 90, 30, 33, 35, 37, 38, 47] explicitly estimate dense motion in the form of optical flow, a dense motion field representing the pixel-wise displacements between two frames, to temporally align input frames and synthesize intermediate frames. Recent advances in optical flow estimation [126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136] have directly propelled the performance of flow-based VFI models. A typical pipeline comprises: (1) estimating either anchor flows ($\mathcal{V}_{0\rightarrow t},\mathcal{V}_{1\rightarrow t}$) or intermediate flows ($\mathcal{V}_{t\rightarrow 0},\mathcal{V}_{t\rightarrow 1}$), (2) applying flow-guided warping [60, 61] of input frames or features ($I_{0},I_{1}$ or $F_{0},F_{1}$), and (3) synthesizing the target frame ($\hat{I}_{t}$) by blending the warped results ($\hat{I}_{0\rightarrow t}/\hat{F}_{0\rightarrow t}$ and $\hat{I}_{1\rightarrow t}/\hat{F}_{1\rightarrow t}$).
The accuracy of the flow critically impacts interpolation quality in this approach, as misalignment directly causes blur and artifacts. Many earlier works [30, 7, 60] adopt off-the-shelf optical flow networks [126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136] to estimate the initial flows. Although these networks offer reliable motion estimation, they are not specifically optimized for the VFI task and often introduce unnecessary architectural complexity. Moreover, they tend to have large model sizes and struggle to handle extreme motions that lie outside the training distribution [125]. To better adapt the motion estimation to the VFI task, a number of methods [16, 5, 6, 74, 62, 35, 59, 81, 97, 83, 55, 54, 53, 125, 57, 100] propose to estimate their own task-oriented flow within their framework, which is optimized jointly with the frame interpolation objective. For example, BiM-VFI [137] distills flow knowledge from an ensemble of flow predictors into a lightweight network tailored for interpolation. GIMM-VFI [88] addresses the noise in flows from pre-trained optical flow estimator (e.g., RAFT [134], FlowFormer [135]) by refining them through a coordinate-based implicit networks. Pseudo ground-truth (GT) strategies are also common, where pseudo GT flow is generated by existing flow networks and used as weak supervision to bootstrap VFI training [16, 79]. These help produce temporally consistent and semantically aligned flows customized for interpolation. With the estimated flow, warping is implemented via either forward [60] or backward [61] warping operation. Depending on the warping operation, it requires different types of flows.
[…]
Figure 6: Comparison of forward and backward warping strategies. (a) Forward warping [60] projects source pixels ($I_{0}$) to their estimated positions in the target frame using $\mathcal{V}_{0\rightarrow t}$. This may introduce holes (unmapped pixels) or collisions (multiple pixels mapped to the same location). (b) Backward warping [61] samples each pixel in the target frame from the source using $\mathcal{V}_{t\rightarrow 0}$. Since sampling is performed at every target location, backward warping naturally produces dense and complete outputs.
Forward warping. Most forward warping-based methods [30, 6, 73, 74, 62, 60, 35, 38, 80, 125, 85] first estimate bidirectional flows $(\mathcal{V}_{0\rightarrow 1},\mathcal{V}_{1\rightarrow 0})$, from which the intermediate flows $(\mathcal{V}_{0\rightarrow t},\mathcal{V}_{1\rightarrow t})$ are linearly interpolated:
$\hat{\mathcal{V}}_{0\rightarrow t}=t\cdot\mathcal{V}_{0\rightarrow 1},\quad\hat{\mathcal{V}}_{1\rightarrow t}=(1-t)\cdot\mathcal{V}_{1\rightarrow 0}.$ (6)
These flows are used to project source pixels to the intermediate frame:
$\hat{I}_{0\rightarrow t}=\overset{\rightarrow}{\mathcal{W}_{f}}(I_{0},\hat{\mathcal{V}}_{0\rightarrow t}),\quad\hat{I}_{1\rightarrow t}=\overset{\rightarrow}{\mathcal{W}_{f}}(I_{1},\hat{\mathcal{V}}_{1\rightarrow t}).$ (7)
However, forward warping introduces structural artifacts such as holes (unmapped regions) and collisions (multiple pixels mapping to the same target position) particularly near motion boundaries [5] as shown in Fig. 6 (a). Unlike backward warping, which ensures dense sampling by mapping each pixel in the target domain, forward warping does not guarantee full coverage due to its source-driven formulation. This intrinsic asymmetry stems from the lack of inverse consistency in optical flow, i.e., $\mathcal{V}_{i\rightarrow j}\neq-\mathcal{V}_{j\rightarrow i}$ in general, particularly under occlusions or non-rigid motion. To address this, SoftSplat [60] proposes a softmax-based splatting mechanism:
$\overset{\rightarrow}{\mathcal{W}_{f}}(I_{0},\mathcal{V}_{0\rightarrow t})=\frac{\vec{\sum}(\exp(Z)\cdot I_{0},\mathcal{V}_{0\rightarrow t})}{\vec{\sum}(\exp(Z),\mathcal{V}_{0\rightarrow t})},$ (8)
where $Z$ denotes a learned importance map (e.g., depth), and the operator $\vec{\sum}$ denotes a differentiable splatting with soft aggregation. The soft aggregation scheme in SoftSplat not only mitigates hole/collision artifacts but also improves the gradient flow by making warping fully differentiable, in contrast to standard splatting operations which are piecewise constant and non-smooth. Despite this, the inherent artifacts make naive forward warping a less favored primary choice.
Backward warping. In contrast, backward warping [5, 33, 76, 81, 59, 55] samples each pixel in the target frame by mapping it back to the input using estimated intermediate flows $(\mathcal{V}_{t\rightarrow 0},\mathcal{V}_{t\rightarrow 1})$. Intermediate flows denote the flows from the unknown target frame to the input frames. Since the target frame is unavailable, it is not straightforward to obtain the intermediate flows. These flows can be approximated via direct prediction [16, 6, 35, 58, 59, 52, 81, 54, 53, 55, 57, 100, 90], flow interpolation [5, 7], or flow reversal [76, 77, 79, 89]. For instance, SuperSloMo [5] employs such linear approximations and further refines them via dedicated subnetworks. The flow interpolation for intermediate flows is defined as:
$\hat{\cal V}_{t\rightarrow 0}$ $=-t\cdot{\cal V}_{0\rightarrow 1}\,\,\textrm{ or }\,\,t\cdot{\cal V}_{1\rightarrow 0}$ (9)
$\hat{\cal V}_{t\rightarrow 1}$ $=(1-t)\cdot{\cal V}_{0\rightarrow 1}\,\,\textrm{ or }\,\,-(1-t)\cdot{\cal V}_{1\rightarrow 0}.$ (10)
To enhance robustness, XVFI [79] introduces Complementary Flow Reversal (CFR), a weighted aggregation strategy that fuses multiple reversed and complementary flows to construct robust intermediate motion fields. This strategy complements the shortcomings of both linear flow approximation and naive flow reversal [76], offering robustness against ambiguities near motion boundaries. Given the intermediate flows, backward warping is applied as:
$\hat{I}_{0\rightarrow t}=\overset{\leftarrow}{\mathcal{W}_{b}}(I_{0},\hat{\mathcal{V}}_{t\rightarrow 0}),\quad\hat{I}_{1\rightarrow t}=\overset{\leftarrow}{\mathcal{W}_{b}}(I_{1},\hat{\mathcal{V}}_{t\rightarrow 1}),$ (11)
where $\overset{\leftarrow}{\mathcal{W}_{b}}$ denotes the backward warping operator [61]. The warped results are blended using occlusion-aware mask $M$ and residual refinement $R$:
$I_{t}=M\odot\hat{I}_{0\rightarrow t}+(1-M)\odot\hat{I}_{1\rightarrow t}+R,$ (12)
The operator $\odot$ denotes element-wise multiplication, or the Hadamard product, which blends the warped frames proportionally based on the occlusion-aware confidence map. Additionally, some methods [5, 76, 75, 79] further incorporates $(1{-}t)$ and $t$ as scalar weights into $M$ to guide time-aware blending. Several methods also exploit auxiliary priors such as depth [7], contextual features [30, 33, 7, 35, 60, 39], or edge information [73, 77, 37] to further guide interpolation. Also, learnable synthesis networks [138] or additional post-processing can further improve sharpness and correct residual artifacts.
Modeling non-linear motion. Many early methods assume linear motion and brightness constancy [16, 5, 30, 33, 7, 73, 76, 74, 6, 35, 60, 85], meaning that objects move along a straight trajectories at constant speed, and pixel intensities remain unchanged. However, these assumptions often fail under real-world scenarios. Quadratic [76, 77, 139, 86] or cubic [62] motion modeling has been proposed to account for acceleration. QVI [76] and EQVI [77] estimate acceleration-aware flows utilizing four input frames. While recent works [137, 90] further explore velocity ambiguity [90], which refers to the ill-posed nature of intermediate motion inference where multiple trajectories yield the same intermediate position, especially under occlusion or acceleration. BimVFI [137] and Zhong et al [90] introduce bidirectional motion fields and time-aware reasoning mechanisms to disambiguate such cases, enabling robust interpolation under occlusion, acceleration, and non-linear motion.
Overall, flow-based methods remain one of the most extensively explored and practically adopted approaches in VFI, owing to their explicit and interpretable modeling of motion trajectories. Their ability to flexibly generate intermediate frames for arbitrary timestamps makes them well-suited for various real-world scenarios requiring variable frame-rate synthesis. Despite these strengths, their performance is sensitive to flow estimation accuracy, particularly under conditions of occlusion, large motion, lighting variation or non-linear motion. As research in optical flow continues to evolve [140, 141], flow-based VFI is expected to further benefit from these developments and remain a foundational component of future VFI approach.
2.2.3 Kernel- and Flow-based Combined
Kernel- and flow-based approaches each offer distinct strengths in VFI. Flow-based methods [16, 30, 7, 33, 74, 35, 37, 38, 97] estimate dense motion fields to align frames in a temporally coherent manner, but their performance degrades due to inaccurate flow estimation or the presence of occlusions. In contrast, kernel-based methods [28, 29, 16, 7, 31, 32, 33, 74, 34, 35, 36, 37, 8, 38, 39, 40, 41, 42, 97, 43, 44, 46, 56, 47] directly synthesize pixels using learned, spatially adaptive convolutional kernels, offering greater robustness in regions with complex motion. However, they are limited by their local receptive field and thus struggle with large displacements.
Hybrid methods combine these complementary approaches by using flow estimation to guide the placement and orientation of learned kernels, yielding both global motion alignment and localized refinement. This combined approach [16, 30, 7, 33, 74, 35, 36, 37, 38, 97] typically begins by estimating optical flows using dedicated flow networks. Some methods adopt off-the-shelf optical flow networks [126, 127, 128, 129, 130, 131, 132, 133, 134, 135, 136] to guide the sampling location or trajectory of adaptive kernels. The kernels are then applied along flow-aligned paths to aggregate motion-aware pixel neighborhoods. MEMC-Net [33] exemplifies this design by integrating PWC-Net for flow estimation and deformable convolution [122] for localized refinement. In this setup, flow fields define the sampling offsets, while the kernel weights are learned to capture residual motion and restore high-frequency content. The predicted flows determine the sampling offsets for each pixel, while the learnable kernels capture residual motion and texture details. More recently, LADDER [47] introduces a lightweight encoder-decoder architecture that jointly estimates motion-aware features and spatially adaptive kernels, reducing complexity while maintaining hybrid modeling capacity.
Despite their accuracy, hybrid approach typically introduces significant computational costs due to the dual pipelines for flow and kernel prediction [41]. To alleviate this, several works [97, 47] adopt encoder-sharing strategies to reduce redundancy and latency. These designs enhance interpolation robustness in scenarios with large displacements, motion ambiguities, or complex occlusion, where single approach-based models often fail. As hybrid architectures continue to evolve, balancing the performance and efficiency remains a central challenge and a promising direction.
2.2.4 Phase-based
An alternative direction in VFI explores the use of phase information to implicitly capture motion cues. In the frequency domain, pixel-wise representations can be decomposed into amplitude and phase components, where temporal phase shifts across frames encode the apparent motion of underlying structures. Phase-based methods [48, 49] exploit this property by estimating motion through local phase variations, rather than relying on explicit correspondence or pixel displacement. To extract and manipulate phase information, most methods adopt multi-scale frequency representations such as complex steerable pyramids [142, 143, 144]. Within this framework, motion is modeled by interpolating both phase and amplitude at each pyramid level. Meyer et al[48] solves this optimization problem explicitly, while later method PhaseNet [49] adopts end-to-end learning strategies.
However, the effectiveness of these methods is fundamentally constrained by the assumption that motion can be approximated as local phase shifts. While this holds for small or moderate motion magnitudes, the assumption breaks down in the presence of large motion, leading to phase ambiguity and aliasing artifacts [145, 146]. As a result, phase-based methods often struggle to preserve fine spatial details or sharp boundaries in high-speed motion scenarios, limiting their applicability in unconstrained, real-world settings. Still, phase representations remain a valuable signal modality and, when combined with other learning-based methods, may help enhance robustness against photometric and structural distortions.
2.2.5 GAN-based
A major limitation of conventional learning-based VFI approaches lies in their reliance on pixel-level loss functions such as $\ell_{1}$, $\ell_{2}$, or perceptual losses based on deep features (e.g., VGG [147]). While these objectives are effective for minimizing reconstruction errors, they often produce perceptually unsatisfying results, characterized by over-smoothed textures and diminished realism [148, 149]. To mitigate this gap, several methods adopt Generative Adversarial Networks (GANs) [150], which demonstrate remarkable performance in synthesizing visually plausible content [151, 152]. These GAN-based VFI methods [93, 101, 103, 74, 36, 94, 96, 95, 102, 97] employ a generator $G$ to synthesize the intermediate frame $\hat{I}_{t}$, and a discriminator $D$ to differentiate between the GT $I_{t}$ and $\hat{I}_{t}$. The generator is optimized using both a reconstruction loss and an adversarial loss, enabling it to produce frames that are structurally coherent with the inputs while exhibiting high perceptual fidelity. Such formulations are particularly effective in hallucinating plausible content in disoccluded regions [153] and enhancing visual details in blurry or textureless areas [74].
Despite their potential, GAN-based models introduce new challenges. They are notoriously difficult to train, often suffering from instability, mode collapse [154], and poor generalization when exposed to motion patterns or scene layouts not well represented in the training data. In such cases, the generator may fail to generalize, leading to artifacts or unrealistic interpolations. As a result, domain adaptation or fine-tuning is often required when applying GAN-based methods to novel environments [155], limiting their scalability in practical deployment.
2.2.6 Transformer-based
Originally proposed for sequence modeling in natural language processing (NLP) [123], the Transformer architecture has been successfully adapted to VFI [50, 51, 52, 53, 54, 55, 56, 124, 113, 57, 117] owing to its strong capacity for capturing long-range dependencies through the attention mechanism [123, 156]. In the context of VFI, where motions often span large spatial and temporal regions with occlusions and deformations, this capability is particularly advantageous. The attention mechanism adaptively weighs features by their relevance to selectively attend to distant yet semantically relevant regions. This is an essential property for synthesizing temporally coherent intermediate frames. The core attention operation is defined as:
$\text{Attn}(Q,K,V)=\text{Softmax}\left(\frac{QK^{\top}}{\sqrt{d}}\right)V,$ (13)
where $Q$, $K$, and $V$ denote the query, key, and value matrices, respectively, and $d$ is the dimensionality of the feature space. This formulation enables the model to focus on spatial-temporal regions that are informative for interpolation, while effectively handling occlusions and appearance changes [37].
Transformer-based VFI methods [50, 51, 52, 53, 54, 55, 56, 124, 113, 57, 117] primarily differ in how they structure attention and encode temporal dependencies. VFIFormer [52] introduces a cross-scale window-based attention (CSWA) mechanism to capture multi-scale dependencies without relying on flow-based motion estimation. Queries are computed from features at the target time, while keys and values are derived from neighboring input frames, enabling direct temporal associations. The multi-scale windowing expands the receptive field, enhancing robustness to complex motion. VFIT [51] employs a hierarchical Transformer operates on multi-resolution features and predicts spatially adaptive blending kernels for fine-grained synthesis. EMA-VFI [53] integrates attention modules with CNNs to reduce overhead, using inter-frame attention to jointly extract motion and appearance cues with improved efficiency.
Despite their effectiveness, this approach suffers from high computational costs. The standard self-attention scales quadratically with the input resolution, posing a bottleneck for HR video inputs. To overcome this, efficient attention designs have been proposed. Swin Transformer [156] reduces complexity via windowed self-attention and shifted windows, while Restormer [157] introduces transposed attention to achieve linear complexity with respect to spatial dimensions. These developments point to a promising direction in which Transformer-based architectures may effectively balance global context modeling with computational efficiency, enabling real-time HR frame interpolation in practical applications.
2.2.7 Mamba-based
Structured State Space Models (SSMs) [98] offer a principled framework for sequence modeling through continuous-time dynamical systems. Among them, Mamba [99] introduces a selective state-space parameterization that combines the recurrent efficiency of recurrent neural networks (RNNs) [158] with the global context modeling capabilities of Transformers. By leveraging linear recurrence and input-dependent gating, Mamba enables long-range dependency modeling with linear complexity. The core of SSM-based modeling is the continuous-time linear time-invariant (LTI) system defined as:
$h^{\prime}(t)$ $=Ah(t)+Bx(t),$ (14)
$y(t)$ $=Ch(t)+Dx(t),$
where $h(t)\in\mathbb{R}^{N}$ is the latent state, $x(t)\in\mathbb{R}$ is the input, and $y(t)\in\mathbb{R}$ is the output. Here, the state size is $N$, with system parameters $A\in\mathbb{R}^{N\times N},\quad B\in\mathbb{R}^{N\times 1},\quad C\in\mathbb{R}^{1\times N},\quad D\in\mathbb{R}$. To incorporate this formulation into deep learning models, the system is typically discretized using the zero-order hold (ZOH) method with a step size $\Delta$. The resulting discrete parameters are computed as:
$\bar{A}$ $=\exp(\Delta A),$ (15)
$\bar{B}$ $=(\Delta A)^{-1}(\exp(\Delta A)-I)\Delta B,$
which yields the following discrete-time recurrence:
$h_{k}$ $=\bar{A}h_{k-1}+\bar{B}x_{k},$ (16)
$y_{k}$ $=Ch_{k}+Dx_{k}.$
which defines the fundamental update rule for state-space sequence modeling.
VFIMamba [100] is the first work to incorporate Mamba into VFI. It introduces a hierarchical architecture based on the S6-based Mixed-SSM Block (MSB) to model temporal dynamics across spatial resolutions. This design enables bidirectional propagation of motion features through structured recurrence, effectively capturing both short- and long-range dependencies. Compared to Transformer-based models [50, 51, 52, 53, 54, 55, 56, 124, 113, 57, 117], VFIMamba achieves lower memory consumption and faster inference while maintaining competitive accuracy, particularly in handling large displacements and preserving high-frequency texture details. MambaFlow [140] presents a Mamba-centric framework for end-to-end optical flow estimation. It extends the modeling capacity of Mamba through two core mechanisms. Self-Mamba captures long-range intra-frame dependencies by applying bidirectional state updates to enrich spatial features with global context. Cross-Mamba, inspired by cross-attention, models inter-frame interactions to improve motion correspondence. Together, these modules collectively improve robustness to occlusion, motion discontinuities, and ambiguous flow regions, making the architecture a promising backbone for VFI. Additionally, other SSM-based models such as MambaIR [159] and MambaIRv2 [160] demonstrate strong performance in image restoration tasks by capturing local detail and global structure with low complexity. The success of these models suggests that structured recurrence offers a compelling alternative to attention mechanisms for spatiotemporal modeling in VFI.
Overall, Mamba provides a promising design space for future VFI frameworks. Its low architectural complexity and reduced sensitivity to hyperparameter tuning offer practical advantages over attention-based designs. However, its ability to handle long-term motion dependencies, occlusion, and non-rigid deformation remains underexplored. Future directions may include exploring hybrid architectures that combine Mamba with local attention or design adaptive recurrence mechanisms conditioned on motion complexity and occlusion patterns.
2.3 Diffusion Model-based
Figure 7: General structure of DM-based VFI framework. The framework receives input keyframes ($I_{0}$, $I_{1}$) and generates intermediate frames ($I_{t}$) through a denoising process. In addition to input keyframes, the model can accept various auxiliary conditioning signals such as images, text, audio, optical flow, or semantic maps via lightweight adapter modules or attention mechanisms.
Traditional deep learning-based VFI methods are predominantly deterministic, assuming a one-to-one mapping between the input frames and the interpolated output [113]. While such models demonstrate strong performance under moderate motion, they encounter fundamental limitations when faced with severe occlusions, significant displacements, or rapid appearance changes, where the underlying motion is ambiguous. In such cases, a deterministic framework cannot fully capture the range of plausible transitions between two frames, often leading to results that diverge from human perceptual expectations [161]. These limitations have prompted the exploration of a generative approach that embraces uncertainty and seeks to synthesize diverse yet semantically coherent interpolations.
DMs [162, 163, 164, 165, 166, 167] have emerged as a powerful generative model, achieving state-of-the-art (SOTA) performance across image [168, 164], video [165, 166], and multimodal generation [169, 170]. Unlike GANs [150] or VAEs [171], which suffer from adversarial instability and posterior collapse respectively, DMs offer stable training, high-fidelity samples, and strong temporal consistency. Inspired by their success in text-to-video (T2V)) [172, 173, 174] and image-to-video (I2V) [175, 169], researchers have recently adapted DMs for VFI [104, 105, 106, 107, 114, 116, 117], expanding the scope of VFI from deterministic interpolation to conditional generative modeling. This paradigm shift redefines VFI as a conditional denoising process, aligning it with the broader concept of Generative Inbetweening [112, 111], which focuses on synthesizing plausible and temporally coherent transitions between sparse keyframes under uncertainty. In this formulation, VFI is modeled as a denoising process conditioned on keyframes, typically denoted as $I_{0}$ and $I_{1}$. Starting from Gaussian noise, a denoising network gradually synthesizes intermediate frames via learned denoising steps. For example, Stable Video Diffusion (SVD) [167] adopts a latent diffusion framework. It first encodes video sequences into compact representations via an encoder $\mathcal{E}(\cdot)$, adds noise in the latent space, and then denoises them using a 3D U-Net [25]. The denoising objective, such as the $\mathbf{v}$-prediction loss [176], encourages accurate reconstruction from noisy inputs:
$\mathcal{L}=\mathbb{E}_{\mathbf{z},\mathbf{c}_{\text{image}},\epsilon,t}\left[\left\|\mathbf{v}-f_{\theta}(\mathbf{z}_{t},\mathbf{c}_{\text{image}},t)\right\|_{2}^{2}\right],$ (17)
where $\mathbf{v}=\alpha_{t}\epsilon-\sigma_{t}\mathbf{z}_{t}$, $\mathbf{z}_{t}$ is the noisy latent at timestep $t$, $\epsilon$ is the GT noise, and $\alpha_{t}$, $\sigma_{t}$ are variance schedule parameters that define the weighting between signal and noise. $\mathbf{c}_{\text{image}}$ denotes the conditioning keyframes.
Initial DM-based VFI models [104, 105] directly predict intermediate content without relying on explicit motion estimation. Building upon this foundation, more recent methods [114, 111] extend VFI into a semantically aware generation task, emphasizing coherent scene evolution over time. To enhance temporal alignment, TRF [112] and ViBiDSampler [115] propose bidirectional sampling and trajectory fusion, facilitating robust interpolation from both forward and backward perspectives without task-specific training. Beyond inference strategies, architectural innovations improve motion control and temporal consistency. EDEN [117] augments the denoising network with a spatio-temporal encoder for global consistency, while MoG [107] integrates motion priors using flow-guided warping in the latent space. As shown in Fig. 7, one of the key strengths of DM-based frameworks lies in their inherent flexibility to incorporate diverse conditioning modalities beyond input keyframes. DMs can seamlessly integrate auxiliary signals such as depth, semantic maps, audio, text, or motion priors via adapter-based [109, 108] or attention-based conditioning pathways. This allows for rich user guidance and semantic control, enabling use cases like such a story-driven animation [108], cross-modal interpolation, and interactive video generation [161]. Framer [161] injects spatial priors into the U-Net via attention mechanisms, while MoG [107] and FCVG [114] adopt ControlNet-like structures [177] to condition the generative process at multiple scales, improving alignment and spatial consistency.
Overall, DM-based VFI provides a new perspective on VFI by decoupling interpolation from deterministic regression and introducing generative modeling as a robust alternative. The ability of DMs to integrate diverse conditioning modalities and to model uncertainty enables flexible and perceptually plausible frame synthesis. However, their computational cost and sampling latency remain notable challenges. Future research directions may include hybrid frameworks that combine explicit motion estimation with generative denoising, as well as curriculum- or cascade-based denoising strategies tailored for HR inputs or temporally long-range interpolation tasks.
3 Learning Paradigm
3.1 Center-Time Frame Interpolation (CTFI)
Figure 8: Comparison of CTFI and ATFI. (a) CTFI only generates a single center-frame at $t{=}0.5$ given two inputs. (b) ATFI can synthesize frames at arbitrary $t\in(0,1)$.
Center-Time Frame Interpolation (CTFI) as shown in Fig. 8 (a), also known as fixed-time interpolation, is a widely adopted learning paradigm in VFI. Here, models are trained on triplets $(I_{0},I_{\frac{1}{2}},I_{1})$ [178, 6, 179, 50, 180], with $I_{0}$ and $I_{1}$ as inputs and $I_{\frac{1}{2}}$ as the GT center-frame. Owing to the simplicity of supervision and precise GT alignment, this paradigm has been dominant in earlier works [92, 28, 29, 16, 76, 50, 36, 59, 51, 113].
Despite its ease of implementation, CTFI suffers from major limitations in real-world scenarios where intermediate frames are required at arbitrary timestamps. Since models are trained to generate only the center-frame at $t{=}\frac{1}{2}$, they inherently lack temporal flexibility for generating frames at other timestamps. For example, to generate a frame at $t{=}\frac{1}{4}$, the model first synthesizes the center-frame, and then recursively generates $\hat{I}_{\frac{1}{4}}$ conditioned on $(I_{0},\hat{I}_{\frac{1}{2}})$. This recursive strategy is inherently sequential, introducing two key drawbacks [5, 79, 181]. First, it increases computational latency and prevents parallel generation, as each intermediate frame depends on the previously synthesized result. Second, it leads to cumulative errors where artifacts in earlier frames propagate through the inference chain, degrading temporal consistency and overall quality. Additionally, CTFI restricts the temporal upsampling factor to powers of two ($2^{n}$), thereby limiting adaptability in diverse frame-rate conversion scenarios such as real-time video streaming or arbitrary slow-motion synthesis.
3.2 Arbitrary-Time Frame Interpolation (ATFI)
In contrast, Arbitrary-Time Frame Interpolation (ATFI) or multi-frame interpolation as shown in Fig. 8 (b), generalizes the task by enabling interpolation at any arbitrary $t{\in}(0,1)$ between two given frames [5, 30, 7, 76, 75, 62, 35, 79, 59, 53, 54, 45, 114, 161]. This paradigm explicitly receives $t$ as input during training and inference, allowing direct synthesis of frames at specified timestamps and supporting continuous-time interpolation. While some earlier methods [7, 73] perform iterative ATFI in a frame-by-frame fashion, such methods often suffer from temporal jitter due to a lack of continuity modeling. In contrast, temporally-aware models [62, 75] predict multiple intermediate frames in one pass, promoting sequence-level coherence and computational efficiency.
Despite its flexibility, ATFI also presents challenges. Training requires HFR datasets to supervise intermediate frames at diverse timestamps. Additionally, ATFI inherently prone to the velocity ambiguity problem, where multiple plausible motion trajectories can lead to the same intermediate position. This often leads models to average over alternatives, resulting in temporal blur. Furthermore, ATFI must account for non-linear motion such as acceleration or abrupt direction changes, phenomena not easily handled under constant-velocity assumptions. These issues are discussed in depth in Sec. 4.4. Despite these challenges, ATFI remains a versatile and powerful paradigm for real-world applications, offering improved flexibility for slow-motion generation, dynamic frame-rate adaptation, and user-controllable playback.
3.3 Training Strategy
3.3.1 CTFI Training Strategy (CTFI-TS)
CTFI-TS builds training triplets $(I_{0},I_{t},I_{1})$ with $I_{t}$ positioned precisely at the center-point between $I_{0}$ and $I_{1}$. These triplets can be generated by sampling three consecutive frames which are uniformly sampled as shown in Fig. 9 (a). This enables the construction of large-scale training datasets without dense manual annotation. During training, models are supervised exclusively at $t{=}0.5$, and no explicit temporal encoding is involved. At inference, the model is similarly evaluated by predicting center-frames at each stride. While efficient, this strategy inherently limits generalization to other timestamps and requires recursive processing for arbitrary-time synthesis.
Figure 9: Comparison of CTFI-TS and ATFI-TS. (a) CTFI-TS samples exactly three uniformly spaced frames per training example, with only the center-frame used as supervision. (b) ATFI-TS uses $(n{+}1)$ uniformly spaced frames from HFR videos, allowing an intermediate frame at arbitrary timestamp $t\in(0,1)$ to serve as supervision target.
3.3.2 ATFI Training Strategy (ATFI-TS)
ATFI-TS constructs training samples from $(n{+}1)$ consecutive frames, using the first and last as inputs $(I_{0},I_{1})$ and the $(n{-}1)$ intermediate frames as supervision targets for their respective times $t\in(0,1)$. Each $t$ is either provided directly or encoded via temporal embeddings [81, 54, 106]. When HFR videos are available, training data can be flexibly constructed by uniformly sub-sampling frames at a desired interval as shown in Fig. 9 (b). As long as the original frame rate of the video is divisible by the desired interpolation factor, any pair of frames can be selected as inputs, and the frames that lie temporally between them can serve as GT supervision targets. This strategy allows models to learn from a wide distribution of motions and time intervals. Inference is fully parallelizable, frames at any $t\in(0,1)$ can be generated independently, making this approach highly efficient and scalable for real-time and high-frame-rate applications. By explicitly modeling time and enabling continuous supervision, ATFI-TS forms the backbone of modern interpolation frameworks seeking generalizability, temporal coherence, and fine-grained control.
3.4 Loss Functions
Loss functions play a critical role in guiding VFI models toward producing temporally coherent and perceptually realistic outputs. They are broadly categorized into reconstruction, perceptual, adversarial, and flow-based losses, each addressing different aspects of the interpolation objective.
3.4.1 Reconstruction Loss
Reconstruction losses supervise the model to minimize the pixel-wise discrepancy between the predicted intermediate frame $\hat{I}_{t}$ and the GT frame $I_{t}^{\text{GT}}$. These losses are typically applied in the RGB space.
$\mathcal{L}_{1}$ Loss is defined as:
$\mathcal{L}_{1}=\left\lVert\hat{I}_{t}-I_{t}^{GT}\right\rVert_{1},$ (18)
which computes the pixel-wise absolute difference between frames.
$\mathcal{L}_{2}$ loss is defined as:
$\mathcal{L}_{2}=\left\lVert\hat{I}_{t}-I_{t}^{GT}\right\rVert_{2}^{2},$ (19)
this loss computes the squared error, yielding smoother gradients but often producing overly smoothed outputs, particularly in high-frequency regions or under motion-induced misalignments [106].
Charbonnier Loss [182] is a differentiable variant of the $\mathcal{L}1$ loss:
$\mathcal{L}_{\text{char}}=\rho(I_{t}^{GT}-\hat{I}_{t}),$ (20)
where $\rho(x)=(x^{2}+\epsilon^{2})^{\alpha}$ is the Charbonnier function, with a small constant $\epsilon$ (typically $10^{-3}$) for numerical stability and $\alpha=0.5$. The loss provides smoother gradients than the $\mathcal{L}_{1}$ loss. Owing to its smooth gradient profile and outlier resilience, Charbonnier loss is frequently adopted in VFI for its balanced sensitivity to both sharp detail and robust training stability.
Laplacian loss [183] compares the Laplacian pyramid decompositions of the interpolated and GT frames to supervise frame synthesis across multiple spatial scales:
$\mathcal{L}_{\text{lap}}=\sum_{i=1}^{l}2^{i-1}\left\lVert L^{i}(\hat{I}_{t})-L^{i}((I_{t}^{GT})\right\rVert_{1},$ (21)
where $L^{i}(\cdot)$ is the $i$-th pyramid level. This encourages alignment of both global structure and fine detail, and is often used in conjunction with $\mathcal{L}_{1}$ loss.
Census loss [184], also referred to as ternary loss, evaluates the structural consistency of local image patches under census transformation [185]. It is defined as:
$L_{\text{cen}}=\psi(I_{t}^{GT},\hat{I}_{t}),$ (22)
where $\psi(\cdot,\cdot)$ is a Hamming-like distance function over census-encoded patches. Due to its robustness against illumination and photometric noise, census loss improves particularly effective in unsupervised or self-supervised VFI frameworks.
3.4.2 Perceptual Loss
To enhance perceptual realism, VFI models often incorporate high-level perceptual losses in addition to pixel-wise criteria. A widely adopted formulation computes feature-level distances using a pre-trained VGG network [23]:
$\mathcal{L}_{\text{per}}=\left\lVert\phi(\hat{I}_{t})-\phi(I_{t}^{GT})\right\rVert_{2}^{2},$ (23)
where $\phi$ denotes the feature extractor. This loss promotes structural consistency and encourages synthesis of semantically aligned textures, especially in challenging visual regions.
3.4.3 Adversarial Loss
To further improve realism, adversarial learning frameworks employ a discriminator $D$ trained to distinguish interpolated frames from real ones. The standard GAN objective is:
$\mathcal{L}_{\text{GAN}}=\mathbb{E}_{I_{t}^{\mathrm{GT}}}[\log D(I_{t}^{\mathrm{GT}})]+\mathbb{E}_{\hat{I}_{t}}[\log(1-D(\hat{I}_{t}))].$ (24)
By optimizing this objective jointly with reconstruction losses, the generator learns to produce sharper and more plausible frames. To enforce temporal consistency, recent works also adopt temporal discriminators, which operate on sequences to distinguish coherent dynamics [74, 62].
3.4.4 Flow Loss
Given that many VFI models rely on motion estimation as an intermediate step, flow supervision becomes critical for improving temporal alignment. Several loss terms are used to regularize or supervise flow prediction.
Smoothness Loss [16] encourages piecewise smooth flow by penalizing abrupt spatial changes:
$\mathcal{L}_{smooth}=\left\lVert\nabla\mathcal{V}_{0\to 1}\right\rVert_{1}+\left\lVert\nabla\mathcal{V}_{1\to 0}\right\rVert_{1}.$ (25)
Warping Loss [5] measures the reconstruction error after warping one frame to the other using estimated flow:
$\mathcal{L}_{warp}=\|I_{0}-\mathcal{W}(I_{1},\mathcal{V})\|_{1}+\|I_{1}-\mathcal{W}(I_{0},\mathcal{V})\|_{1},$ (26)
where $\mathcal{W}$ denotes the warping operator.
First-order Edge-aware Smoothness Loss [79] is designed to preserve sharp motion discontinuities, this loss attenuates regularization near edges:
$\mathcal{L}_{edge}=\sum_{i=0,1}\exp\left(-e^{2}\sum_{c}|\nabla_{x}I^{0}_{tc}|\right)^{\top}\cdot|\nabla_{x}\mathcal{V}^{0}_{ti}|,$ (27)
where edge strengths are computed via image gradients and used to modulate the smoothness penalty.
4 VFI Challenges
Figure 10: Representative challenges in VFI. (a) Large motion makes it difficult to establish accurate correspondences across frames, especially in cases involving fast-moving objects, deformable structures, or significant camera motion. (b) Occlusion introduces ambiguity, as some regions in the intermediate frame are not visible in either of the input frames, making it unclear what content should be synthesized. (c) Lighting variations, such as shadows, reflections, or changes in illumination, violate brightness constancy assumptions and hinder accurate motion estimation. (d) Non-linear motion refers to changes in motion speed or direction over time, making it difficult to infer intermediate positions.
Despite extensive progress in VFI, several representative challenges consistently remain difficult across approaches, limiting real-world performance. As shown in Fig. 10, these include large motion [60, 58], occlusion [87], lighting variation, and non-linear motion [90, 137].
4.1 Large Motion
Large motion refers to scenarios where objects undergo substantial displacement between consecutive frames. As shown in Fig. 10 (a), this includes articulated movements (e.g., a person leaning left to right) or abrupt camera motion, which result in wide spatial shifts across the image plane. Such motion is prevalent in real-world videos and presents a fundamental challenge in VFI due to the difficulty of establishing accurate correspondences over long spatial ranges.
To accurately synthesize an intermediate frame, the model must identify where each pixel from the first frame $(I_{0})$ has moved in the following frame $(I_{1})$ which denotes motion or correspondence estimation. When the motion is small, this is relatively straightforward because corresponding pixels remain close. However, large motion induces long-range dependencies that exceed the receptive field of standard networks. Moreover, appearance changes and occlusions further hinder accurate estimation by introducing discontinuities in motion and visibility. To address this, many VFI models adopt a coarse-to-fine hierarchical framework, where large displacements are estimated at low-resolution (LR) feature maps and progressively refined at higher resolutions. RIFE [59] employs multi-scale residual flow refinement, enabling robust alignment under wide motion ranges. FILM [83] leverages a feature pyramid for flow estimation and lightweight synthesis, explicitly targeting fast motion and blur scenarios. Similarly, IFRNet [81] improves motion encoding through a motion-aware feature extractor and an intermediate flow refinement block. In addition to these designs, some models further enhance alignment under large displacements by leveraging bidirectional motion modeling [5, 58, 56, 57] or attention mechanisms [97, 55, 161]. ABME [58] proposes asymmetric bilateral estimation, predicting forward and backward flows independently to improve robustness under occlusion. BiFormer [55] incorporates deformable attention across bidirectional contexts, enabling the model to dynamically attend to semantically relevant but spatially distant regions, an effective strategy for capturing non-local motion patterns.
Despite architectural differences, these models all share a common objective of expanding the receptive field effectively while maintaining spatial precision. To this end, many methods combine multi-scale refinement, attention-based global matching, and motion-aware modules, enabling them to handle wide-range motion more effectively. Such designs have demonstrated strong performance on benchmarks involving extreme motion, most notably X4K1000FPS [79], which provides 4K videos at 1000fps along with dense GT for fine-grained evaluation. Following the introduction of X4K1000FPS, several HR datasets [186, 187] have been proposed to further benchmark performance under high-speed and large-displacement conditions. By providing more realistic and challenging settings, these datasets have enabled better training and evaluation of VFI models in unconstrained environments. As a result, the availability of such benchmarks has accelerated the development of more robust architectures capable of preserving motion detail and fidelity under large displacements.
4.2 Occlusion
Achieving high-quality (HQ) interpolation demands accurate motion estimation as well as a proper understanding of occlusions. Otherwise, severe artifacts are likely to appear in the interpolated frames, particularly near motion boundaries. For two consecutive input frames, certain pixels in the intermediate frame may not correspond to any observable region in either input, creating ambiguity in determining the correct content for these occluded regions [188]. As shown in Fig. 10 (b), such occlusions can occur when previously hidden objects become visible or when objects move toward the camera, revealing regions that were not seen in either input. Naively blending warped inputs often results in severe artifacts, most notably ghosting artifacts [87], where an object is not only incorrectly projected from its previous location but also appears as a duplicate at its correct position due to the lack of sufficient visual cues. This is especially problematic in disoccluded regions, areas newly revealed in the intermediate frame but absent in both inputs, such as when an object emerges from behind another or moves directly toward the viewpoint. In these cases, the absence of visual evidence introduces ambiguity, making it unclear what content should be synthesized. To resolve this, modern VFI methods incorporate explicit occlusion reasoning to guide the synthesis process.
A common approach involves estimating soft occlusion masks that weight the pixel contributions from each frame [5, 74, 7, 33, 6]. SuperSloMo [5] jointly predicts bidirectional flow and occlusion masks to exclude unreliable pixels during frame blending. SoftSplat [60] improves upon this by introducing a differentiable softmax visibility map that enables confidence-weighted forward warping. OCAI [87] further incorporates forward-backward consistency checks [184] to identify unreliable flow regions and applies targeted masking and flow inpainting to recover missing structures. In addition to visibility maps, auxiliary cues such as context and depth also improve occlusion handling. CtxSyn [30] integrates warped context features alongside frames to guide synthesis with spatial awareness. DAIN [7] estimates occlusion areas using depth information and leverages neighboring contextual cues to fill the missing regions.
Overall, occlusion-aware VFI remains a critical challenge, particularly in dynamic scenes with depth discontinuities or disoccluded motion. As such, SOTA models increasingly combine multiple strategies, such as masking, depth priors, feature similarity, or forward-backward consistency [184, 87] to recover plausible content in ambiguous regions and maintain temporal coherence in the output.
4.3 Lighting Variation
Lighting variation refers to temporal changes in illumination, shadows, reflections, or exposure across consecutive frames as shown in Fig. 10 (c). These variations can significantly degrade the quality of interpolation, as they violate the basic assumption of brightness constancy [27, 189], which is widely adopted in many optical flow and motion estimation methods. This assumption presumes that the intensity of a surface patch remains constant across time as it moves, allowing pixel-wise correspondences to be inferred from photometric similarity. However, in practice, lighting changes can cause the same object to appear drastically different between frames, resulting in erroneous motion estimation and visually inconsistent interpolations.
To mitigate this, alternative representations have been proposed. Phase-based methods [48, 49] operate in the frequency domain, where motion is encoded as phase shifts rather than intensity differences. These models leverage phase information that remains stable under lighting fluctuations, yielding temporally coherent interpolations even in the presence of flickering or exposure variation. More recently, Transformer-based architectures have shown robustness to photometric inconsistencies. TTVFI [124] aligns motion features across temporal trajectories using attention, enabling the model to blend semantically aligned tokens rather than relying on raw pixel intensities. This higher-level representation effectively helps suppress errors induced from inconsistent lighting, producing perceptually coherent results.
Although lighting variation has received less attention than large motion or occlusion problem, existing methods suggest that photometric-invariant features, frequency-domain modeling, and attention-based alignment provide viable solutions. Continued exploration of these strategies could further enhance the robustness of VFI models in unconstrained environments.
4.4 Non-linear Motion
Many early VFI methods [16, 5, 30, 33, 7, 73, 76, 74, 6, 35, 60, 85] assume linear or uniform motion between input frames. Under this assumption, objects move along straight trajectories at constant velocity, allowing motion estimation based on simple temporal interpolation. Flow-based [16, 5, 33, 6], kernel-based [33], and even phase-based models [49] often rely on this assumption implicitly. However, in real-world scenarios, motion is frequently non-linear due to acceleration, deceleration, or directional change. As shown in Fig. 10 (d), a sliding ball accelerates along a curved path, violating the linear motion prior and introducing significant estimation error.
To address these limitations, researchers have proposed higher-order motion modeling that extends beyond linear assumptions. Since most existing methods operate on only two input frames, they are inherently under-constrained and forced to assume simple motion. To overcome this, several methods incorporate multiple input frames (typically four) to capture richer temporal variations and better approximate non-linear motion. QVI [76] introduces a quadratic motion model that fits second-order trajectories over four consecutive input frames. Specifically, it takes $(I_{-1},I_{0},I_{1},I_{2})$ as inputs and predicts an intermediate frame $I_{t}$ for arbitrary $t\in(0,1)$. By modeling both velocity and acceleration from surrounding frames, QVI enables the network to better handle curved or time-varying motion paths. This parametric formulation allows the model to explicitly account for motion curvature. EQVI [77] further refines by combining offset-based warping with temporal embeddings, improving precision and robustness under complex motions. More recently, IQ-VFI [86] introduces an implicit motion representation using a coordinate-based MLP that adapts to arbitrary motion patterns without requiring predefined trajectory assumptions. These works collectively emphasizes the importance of modeling non-linear motion directly, especially in multi-frame settings. However, these simple mathematical models cannot completely capture the complexities and irregularities of real-world motions.
As the field progresses, a new challenge velocity ambiguity [90] appears. When only two frames are available, multiple plausible motion trajectories can explain the observed displacement, making the underlying motion inherently under-constrained. This ambiguity becomes especially pronounced in scenes involving curved motion or directional switches, such as bouncing balls or rotating limbs. To tackle this, Zhong et al. [90] introduces a velocity embedding module that learns to disambiguate temporal dynamics by jointly reasoning over motion direction and temporal consistency. It separates appearance modeling from motion estimation, which enhances robustness in complex scenes. BiM-VFI [137] takes a complementary perspective by designing an explicit bidirectional motion descriptor. Its Bidirectional Motion Fields (BiM) encode angular and magnitude differences relative to the intermediate time, enabling accurate modeling of curved, asymmetric, or velocity-changing trajectories. BiM-VFI further integrates these representations into a BiM-guided flow estimator and motion-aware refinement network, yielding temporally coherent results in non-linear regimes. These recent advances signal a shift from rigid linear motion priors toward flexible, context-aware motion modeling. By extending temporal supervision and refining motion representations, either through quadratic formulations, implicit embeddings, or directional velocity fields, modern VFI methods now offer significantly improved performance in complex motion scenarios that were previously underexplored.
5 Datasets and Evaluation
5.1 Datasets
To facilitate training and evaluation across varying temporal resolutions and motion complexities, numerous VFI datasets have been developed. Table I provides a high-level summary of commonly used datasets categorized into triplet and multi-frame types. We describe each dataset in detail below.
5.1.1 Triplet Datasets
Early learning-based VFI approaches primarily rely on triplet datasets, where two input frames are used to predict the temporally centered GT frame. This configuration aligns with CTFI settings (Sec. 3.1). Some datasets are further extended to seven-frame sequences [6] for evaluating frame-rate upsampling.
  • Middlebury [189]: Originally designed for optical flow, Middlebury contains short video clips with moderate complexity. Its small size limits scalability, but it remains a standard benchmark for consistency evaluation.
  • UCF101 [178, 16]: A human action dataset from which a small subset of triplets is used for VFI. Due to its LR and simple motion, it is mainly used for training or sanity checks.
  • Vimeo90K [6]: A widely adopted benchmark with diverse scenes and consistent format. It offers clean supervision and balanced motion complexity, making it ideal for comparative analysis.
  • SNU-FILM [50]: Constructed from high-speed footage and categorized by motion difficulty, SNU-FILM enables evaluation across varying levels of motion, occlusion, and blur.
  • ATD-12K [180]: A large-scale animation dataset with rich stylistic diversity. Its variation in artistic textures and motion patterns supports both general-purpose and domain-specific evaluation.
5.1.2 Multi-frame Datasets
TABLE I: Summary and comparison of popular datasets for VFI.
The dataset types
T
represents Triplet dataset,
M
represents Multi-frame dataset.
Dataset Venue Type Resolution Split #Videos / #Triplets URL
[…]
test 12
[…]
test 379
[…]
test 3,782
[…]
test 1,240
[…]
test 2,000
[…]
test 8
[…]
test 195
[…]
test 20
[…]
test 11
[…]
test 12
[…]
test 10
[…]
test 11
[…]
test 15
[…]
test -
[…]
test 53,494
[…]
test -
Multi-frame datasets enable dense temporal supervision and are commonly used in both CTFI and ATFI (Sec. 3.2) settings. They support flexible frame sampling and facilitate evaluation under diverse temporal intervals.
  • Xiph [195, 60]: A curated set of 4K video sequences designed for assessing interpolation fidelity in subtle motion settings.
  • KITTI [190]: Captured in autonomous driving scenarios, KITTI poses unique challenges with sparse GT and large ego-motion.
  • Sintel [192]: A synthetic dataset rendered from the Sintel film, offering photorealistic motion and structured flow annotations.
  • DAVIS [191]: Originally for segmentation, DAVIS features complex object motion, occlusion, and deformation, offering rich dynamics for interpolation.
  • Adobe240 [179]: Collected at 240fps, this dataset captures real-world motion blur and lighting changes, ideal for fine-grained temporal modeling.
  • GOPRO [193]: Featuring high-frame-rate recordings with handheld cameras, GOPRO provides realistic non-linear motion and defocus blur.
  • HD [33]: A subset of HR content from Xiph, with sharper motion content suited for realistic evaluation.
  • X4K1000FPS [79]: A premier benchmark for ultra-slow motion and long-range interpolation, thanks to its dense 1000fps and 4K capture settings.
  • WebVid-10M [194]: A large-scale web video corpus originally built for text-video tasks. Its size and diversity support generative VFI when properly filtered.
  • LAVIB [187]: Designed for large-scale, diverse-domain evaluation with balanced splits and curated subsets for out-of-distribution testing.
  • OpenVid [186]: A text-video dataset supporting multi-modal VFI and DM-based interpolation research via dense, aligned samples.
5.2 Data Augmentation
Modern VFI models incorporate spatial and temporal data augmentation to improve generalization and prevent overfitting. A widely adopted strategy is patch-based cropping, where fixed-size patches (e.g., $128\times 128$ or $256\times 256$) are randomly extracted from HR inputs [29, 62, 81, 54]. This not only reduces memory and computational costs but also encourages localized motion learning while mitigating spatial overfitting to scene layout or object positioning. Furthermore, random cropping prevents the model from overfitting to spatial priors such as background layout or object location, thereby improving robustness across spatial contexts [29]. Additional spatial augmentations, such as horizontal/vertical flipping and random rotation, enhance appearance diversity and promote invariance to orientation and perspective changes. These augmentations enable the model to remain invariant to directional biases and better generalize to unseen spatial transformations.
Temporal augmentation is equally critical in sequential modeling. Frame order reversal [81, 54] is commonly applied, wherein sequences like $(I_{0},I_{1},I_{2})$ are reversed to $(I_{2},I_{1},I_{0})$. In CTFI, this augmentation preserves the center-frame $I_{1}$ while exposing the model to symmetric motion trajectories [5, 50]. Similarly, in ATFI settings, reversing sequences ensures temporal consistency under bidirectional motion. For example as shown in Fig. 9 (b), consider an input triplet $(I_{0},I_{\frac{1}{3}},I_{1})$ used to supervise interpolation at $t{=}\frac{1}{3}$. By reversing the sequence to $(I_{1},I_{\frac{1}{3}},I_{0})$, the relative time becomes $(1{-}\frac{1}{3}){=}\frac{2}{3}$. This simple yet effective strategy enables the model to learn temporally symmetric representations, thereby improving generalization across motion directions and enhancing robustness in bidirectional synthesis.
Overall, these augmentation act as effective regularizers, enabling VFI models to generalize across diverse motion scales, temporal patterns, and visual variations. Integrating these schemes has become a foundational component of both CTFI and ATFI training pipelines.
TABLE II: Summary of evaluation metrics for VFI.
[…]
Category Metric Interpolation Quality Reference Frame
[…]
5.3 Evaluation Metrics
To facilitate comprehensive assessment of VFI models, various metrics have been proposed to capture different aspects of visual quality and temporal coherence. Table II summarizes commonly used evaluation metrics categorized into image-level, perceptual, and video-level types.
5.3.1 Image-level Metrics
Image-level metrics assess the quality of individual interpolated frames with respect to GT references. These pixel-centric evaluations focus on spatial accuracy without considering temporal dependencies across video sequences.
Peak Signal-to-Noise Ratio (PSNR) quantifies reconstruction fidelity based on the mean squared error (MSE) between interpolated frame and GT frame. While higher PSNR reflects better numerical similarity, it often fails to align with human perception, especially for high-frequency or perceptually salient regions.
Structural Similarity Index (SSIM) [196] evaluates local structural integrity by comparing luminance, contrast, and texture patterns. SSIM values range in $[-1,1]$, with higher values indicating stronger structural alignment. Though more perceptually aligned than PSNR, SSIM may still overrate visually implausible outputs if global structure is preserved.
Interpolation Error (IE) [189] computes the root-mean-square error (RMSE) between interpolated frame and the GT frame. Despite being intuitive, IE shares limitations with PSNR in terms of perceptual relevance.
5.3.2 Perceptual Metrics
Perceptual metrics aim to assess the semantic plausibility, texture fidelity, and structural realism of interpolated frames, often aligning better with human visual preferences.
Natural Image Quality Evaluator (NIQE) [197] is a no-reference score derived from deviations to natural image statistics. Lower values reflect more natural, HQ frames.
Fréchet Inception Distance (FID) [198] measures the Fréchet distance between the feature distributions of generated frames and GT frames using a pre-trained Inception network [208]. Lower FID indicates better semantic alignment.
Learned Perceptual Image Patch Similarity (LPIPS) [199] measures perceptual similarity using deep features from pretrained networks. It is robust to minor misalignment and sensitive to semantic differences. Lower LPIPS signifies better perceptual quality.
FloLPIPS [200] extends LPIPS by applying motion-aware weighting based on optical flow. It emphasizes visual fidelity in regions undergoing large displacement.
STLPIPS [201] improves LPIPS by incorporating shift-tolerant feature matching, enhancing robustness to slight misalignments.
DISTS (Deep Image Structure and Texture Similarity) [202] separately evaluates texture and structure similarity using deep features. It balances local detail and global consistency.
5.3.3 Video-level Metrics
These metrics assess spatiotemporal coherence over video sequences, which is essential for realistic and temporally consistent interpolation.
VSFA [203] is a no-reference model trained on human labels. It estimates perceptual quality by aggregating deep features with a recurrent network. Lower scores suggest better perceived video quality.
tOF [204] computes temporal optical flow consistency across frames. Lower tOF values indicate smoother motion continuity.
Fréchet Video Distance (FVD) [205] measures the Fréchet distance between distributions of deep features extracted from real and generated videos using a pre-trained Inflated 3D ConvNet (I3D) [209]. Lower FVD values reflect stronger temporal and perceptual realism.
Fréchet Video Motion Distance (FVMD) [206] improves upon FVD by disentangling motion and appearance, focusing more explicitly on dynamic consistency.
VBench [207] is a multi-dimensional benchmark that scores video models across motion fidelity, coherence, and realism. It enables large-scale reference-free evaluation using semantic video representations.
6 Applications
6.1 Event-based VFI
Event-based Video Frame Interpolation (EVFI) [210, 211, 212, 213, 214, 215, 216, 217, 218, 219, 220, 221, 222, 223, 14] aims to improve interpolation accuracy by leveraging the unique advantages of event cameras. Unlike conventional frame-based cameras that capture full images at fixed intervals, event cameras [224], which are bio-inspired vision sensors [225], asynchronously record per-pixel brightness changes, referred to as “events”, triggered when a contrast threshold is exceeded. These sensors offer key benefits such as ultra-high temporal resolution, high dynamic range, and low latency, making them ideal for scenarios involving rapid motion or challenging lighting. Consequently, event cameras have gained traction in VFI research, especially where traditional RGB frames suffer from motion blur or low temporal fidelity [212, 216, 218].
One of the early models, TimeLens [212], estimates optical flow directly from event streams and synthesizes intermediate frames accordingly. Later models such as TimeReplayer [216] and EGVD [223] improve performance by jointly estimating motion and appearance. TimeLens-XL [221] enhances any-time interpolation capability by optimizing flow and frame synthesis iteratively. Despite these advances, EVFI models remain sensitive to synthesis errors, as inaccuracies can accumulate over time, leading to temporal artifacts.
Despite their strengths, EVFI models face practical challenges. Capturing real event streams requires specialized neuromorphic sensors, which are often expensive and less accessible than conventional cameras. Moreover, collecting large-scale event datasets with dense GT labels is especially challenging due to the asynchronous nature of event recordings. As a result, several studies [226, 227, 228, 229] exploit event simulation from standard camera, simulating the event stream from continuous images or video sequences. Kaiser et al. [226] simulates positive or negative events by thresholding the intensity change between consecutive frames. Pix2NVS [227] estimates per-pixel luminance from video to synthesize event-like representations, aligned to frame intervals.
6.2 Cartoon VFI
Producing traditional 2D animation is labor-intensive [230], requiring artists to manually draw multiple in-between frames. VFI offers a means of automating this process by generating plausible intermediate frames, thereby reducing production time and cost [230, 231].
However, cartoon videos exhibit distinct characteristics compared to real-domain videos: they feature exaggerated motion, minimal texture, flat color regions, and sharp contours, which pose challenges to correspondence-based methods. To address this, domain-specific models have been proposed [180, 232, 233, 109, 234, 235]. Notably, ToonCrafter [109] adopts a generative framework rather than relying on explicit motion estimation. Recent efforts aim to build models that generalize across both cartoon and real domains by leveraging diverse training data or domain adaptation techniques [114, 161, 116].
A major bottleneck in cartoon VFI research is the absence of standardized, HQ datasets. While ATD-12K [180] provides a useful benchmark, its triplet-only format restricts its utility in ATFI settings. As a result, future progress will depend on the release of open, multi-frame cartoon datasets that enable fair and reproducible evaluation.
6.3 Medical Image VFI
VFI is also increasingly applied in medical imaging to reconstruct temporally dense 4D sequences from sparsely acquired volumetric scans [236, 237, 238]. Modalities like CT and MRI face acquisition constraints due to radiation exposure and long scanning times [237], leading to coarse temporal sampling. VFI offers a means to generate intermediate volumes that enhance temporal resolution without incurring additional scan overhead. Medical VFI models must account for subtle anatomical motions and preserve fine structural detail critical for clinical interpretation. CPT-Interp [238] uses continuous motion field modeling, while DU4D [237] proposes an unsupervised interpolation framework that does not rely on GT annotations. These approaches enhance applicability in settings where labeled 4D medical datasets are scarce. Nonetheless, challenges remain. Ensuring clinical validity, minimizing hallucinated content, and establishing domain-specific evaluation metrics are ongoing concerns. Future research will likely explore physiology-aware modeling, uncertainty quantification, and benchmark design specific to 4D medical imaging tasks.
6.4 Joint Task
Recent studies have explored jointly performing VFI with other LLV tasks such as super-resolution (SR) [239, 240, 241, 8, 242, 243] and deblurring [244, 245, 246, 181, 247, 248]. Such joint formulations exploit the inherent correlation between spatial and temporal cues in video sequences [240]. For instance, space-time video super-resolution (STVSR) jointly upsamples resolution and frame rate by leveraging spatial details to enhance motion estimation and vice versa [240]. Shared representations enable efficient feature reuse, reduce redundancy, and facilitate joint optimization. Models such as FISR [240] and MOTIF [243] exemplify this integrated approach. Joint deblurring and interpolation addresses scenarios involving both motion blur and low frame rates. Instead of applying deblurring followed by VFI in a cascade, end-to-end models [139, 244, 245, 246, 181, 247, 248] simultaneously estimate clean and interpolated frames, resulting in improved temporal consistency and visual clarity. These multitask designs improve robustness and efficiency, particularly under real-world degradation, and suggest promising directions for unified LLV modeling.
7 Future Research Directions
7.1 Video Streaming Service
The widespread adoption of real-time video services, including video conferencing and adaptive streaming, presents a growing need for bandwidth-aware video delivery under constrained networks [249]. VFI offers a promising solution by enabling keyframe-only transmission while synthesizing intermediate frames on the client side, thus maintaining visual fluidity at lower bitrates. While early methods confirm its potential for rate reduction, practical deployment remains scarce due to model complexity, inference latency, and platform limitations. Future directions include the development of ultra-lightweight architectures that can operate on mobile or edge devices with limited compute resources. Moreover, adaptive interpolation strategies that jointly consider network bandwidth, scene motion complexity, and perceptual saliency are needed. Learning-based rate control, where the interpolation fidelity is dynamically modulated, could enable bitrate–quality trade-offs tuned in real-time. Joint optimization pipelines that integrate VFI models into codecs or reinforcement learning-based streaming agents may unlock robust low-latency video systems. In particular, methods that unify frame interpolation with residual-based encoding and decoding schemes could blur the boundary between generation and compression, laying the groundwork for next-generation streaming protocols.
7.2 All-in-One LLV Video Restoration
While all-in-one models have shown promising results in image restoration [250, 251, 252], equivalent progress in the video domain particularly in unified LLV frameworks, remains limited. Current LLV restoration pipelines remain fragmented, with VFI, denoising, deblurring, and SR often treated as separate tasks. This modularity, while convenient for controlled benchmarking, limits model robustness under real-world degradations that involve complex mixtures of temporal and spatial artifacts. A promising direction is the development of unified, all-in-one architectures that perform multiple LLV tasks jointly, where VFI is not treated as a standalone module but as an integral part of a broader restoration framework. The interpolated frames can offer temporally consistent guidance for denoising or deblurring, while super-resolved outputs can enhance motion estimation accuracy. Cross-task consistency losses or multi-task learning objectives can foster synergistic improvements. Moreover, transformer- or diffusion-based architectures with spatio-temporal attention mechanisms are naturally suited to this multi-task paradigm, as they can encode long-range dependencies and modulate task-specific pathways via conditioning.
7.3 3D and 4D Scene Understanding
VFI research remains largely grounded in 2D image space, often assuming planar motion and flat appearance fields. However, the increasing prevalence of 3D-aware applications in AR/VR, robotics, and multiview rendering calls for VFI methods that explicitly account for the underlying geometry of dynamic scenes. Recent works in 4D scene modeling using temporal neural fields [253], dynamic Gaussians [254], and neural point representations [255] suggest that temporally coherent synthesis is possible when motion is modeled in 3D space. Integrating VFI into such pipelines enables physically plausible interpolation that respects depth, occlusion, and parallax. Depth-conditioned flow, pose-aware synthesis, or geometry-aware latent spaces may serve as intermediate representations. Furthermore, VFI models can be extended to generate novel viewpoints, enabling geometry-consistent interpolation across spatial and temporal domains. Applications span from time-synchronized multiview interpolation to immersive scene reconstruction from sparse video inputs. Future work may explore co-training paradigms that fuse interpolation and view synthesis losses, jointly supervising geometry, appearance, and motion fields across space-time volumes.